OpenStudy (jiteshmeghwal9):

Find out \(x\) and \(y\) :- \(3\dfrac{7}{x} \times y\dfrac{3}{15}=8\)

Parth (parthkohli):

Diophantine equation?

OpenStudy (jiteshmeghwal9):

o_O

Parth (parthkohli):

\[\dfrac{3x + 7}{x} + \dfrac{5y + 1}{5} = 8\]

OpenStudy (jiteshmeghwal9):

Diophantine equation=? It's the question of N.T.S.E algebra, i don't know nothing more than this

Parth (parthkohli):

Do you need just the integer solutions?

OpenStudy (jiteshmeghwal9):

\[\frac{15x+35+5yx+x}{5x}=8\]\[16x+35+5yx=40x\]\[35+5yx=24x\]& yes @ParthKohli i only need an integer solution :)

Parth (parthkohli):

Hmm... complete the square?

OpenStudy (experimentx):

35 = (24 - 5y)x try factorizing 35

OpenStudy (anonymous):

3 + 7/x + y+1/5 =8 7/x + y = 24/5 7/x = (24-5y)/5 So, x=5n and 7=(24-5y)n y=4 + (4n-7)/5n

OpenStudy (anonymous):

Since 5n>4n-7 for n>0 No positive integer solution

OpenStudy (jiteshmeghwal9):

ans. 11, 2

OpenStudy (experimentx):

looks like you didn't do the algebra correctly .. http://www.wolframalpha.com/input/?i=integer+solution+35%2B5yx%3D24x

OpenStudy (jiteshmeghwal9):

Hey! parth i found a page on diophantine equations

Parth (parthkohli):

Yes! @experimentX I was searching for the factoring trick =)

OpenStudy (experimentx):

this is a classic trick 35 = 1x35 35 = 7x5 -------------- 1*35 = x * (5y - 24) compare these types of factors and see if you get integer values. just an example :)

Parth (parthkohli):

@jiteshmeghwal9 Search for Art of Problem Solving's video: Simon's Favorite Factoring Trick

OpenStudy (jiteshmeghwal9):

Then is the answer given in my book is wrong ?

OpenStudy (experimentx):

no ... what i copied was erroneous data

OpenStudy (jiteshmeghwal9):

ohh..k

OpenStudy (experimentx):

hah!! how could this yield same result http://www.wolframalpha.com/input/?i=integer+solution+7%2Fx+%3D+%2824-5y%29%2F5

OpenStudy (experimentx):

perhaps the algebra was right and the answer on book is wrong.

OpenStudy (anonymous):

(3 + 7/x) (y +3/15) =8 (y+3/15) = 8/(3 +7/x) (15y+3)/15 = 8x/(3x+7) => 8x =(15y+3)n =>(3x+7)=15n

OpenStudy (experimentx):

even the solution is pretty clear from this equation 7 (24-5y) -- = -------- x 5

OpenStudy (jiteshmeghwal9):

24-5y=7 5y=17 y=17/5 x=5

OpenStudy (jiteshmeghwal9):

Is this the right answer ?

OpenStudy (jiteshmeghwal9):

@experimentX ?

Parth (parthkohli):

But \(17/5\) is not an integer :P

OpenStudy (experimentx):

woops!! sorry ... i guess I made error there. Let me check again. Don't have much time.

OpenStudy (experimentx):

3 + 7/x + y +1/5 =8 (3x+7)/x = 8 - y - 1/5 -------------------- \[ \frac{3x+7}{x} = \frac{5(8-y)-1 }{5} \]

OpenStudy (experimentx):

multiply both by 'k' ... numerator and denominator, and look for integer values. x = 5k 15k + 7 = 5(8-y)-1

OpenStudy (sirm3d):

the original problem is \[3\frac{7}{x}\times y\frac{3}{15}=8\] but your beginning line is \[3\frac{7}{x}\color{red} + y\frac{3}{15}=8\] so you probably will not get \(x=11,y=2\)

OpenStudy (jiteshmeghwal9):

I'm totally confused

OpenStudy (jiteshmeghwal9):

LOL, @sirm3d u wrote we'll \(\LARGE\star not\) get x=11, y=2, haha :)

OpenStudy (anonymous):

(3 + 7/x) (y +3/15) =8 (y+3/15) = 8/(3 +7/x) (15y+3)/15 = 8x/(3x+7) => 8x =(15y+3)n =>(3x+7)=15n

OpenStudy (anonymous):

I guess NO SOLUTION

OpenStudy (jiteshmeghwal9):

Everyone has their own opinion, what's the real answer ???

OpenStudy (sirm3d):

i found the solution \(11,\;y=2\)

OpenStudy (anonymous):

x =(7y+21)/(111-3y)

OpenStudy (jiteshmeghwal9):

7(y+3)/3(37-y)

OpenStudy (anonymous):

If solution exists 7y+21>=111-3y =>y>=9

OpenStudy (sirm3d):

\[\vdots\\(3x+7)(25y+1)=40x\\40x=(1,40x),(2,20x),\dots (40,x);(x,40),(2x,20),\dots,(40x,1)\]

OpenStudy (sirm3d):

typo error above. it should be \(5y+1\) by choosing \((40,x)\) as factors of 40, \[3x+7=40,\;5y+1=x\]

OpenStudy (whpalmer4):

x=21,y=40 is a solution for the problem as initially stated

OpenStudy (sirm3d):

\[3x+7=40\Rightarrow x=11\\5y+1=x=11\Rightarrow y=2\]

OpenStudy (anonymous):

(3 +7/x) (y+3/15) =8 . . . y=(31x+21)/(15x+35) y=(30x+70+x-49)/(15x+35) y=2 + (x-49)/(15x+35)

OpenStudy (anonymous):

when x=-6 y = 2 + (-6-49) / (15 *(-6) +35) y=3

OpenStudy (anonymous):

So, x=-6 and y=3

OpenStudy (anonymous):

Now, Since I15x+35I > Ix-49I for x<=-7 and x>=1 x cannot be l than and equal to -7 and cannot be greater than 1

OpenStudy (anonymous):

And NO SOLUTION in the interval -5<=x<=0

OpenStudy (anonymous):

So, If I did my algebra correct the only solution is x=-6 and y=3

OpenStudy (whpalmer4):

And which equation do you find x=-6 and y = 3 to solve?

Parth (parthkohli):

Achcha toh upar waali equation ko simplify karke kya nikalta hai?

OpenStudy (jiteshmeghwal9):

35+5yx=24x

Parth (parthkohli):

Toh hum diophantine equations mein factorization kar sakte hain jinka form hai \(axy + bx + cy =0\) ya phir \(axy + bx + c\) ya phir \(axy + by + c\).

Parth (parthkohli):

Trick hai ki humein woh constant ko ek side laana hai aur left side factor karna hai

OpenStudy (jiteshmeghwal9):

accha thik hai toh equation 35+5yx-24x=0 mein change ho jaayegi

Parth (parthkohli):

Ab equation mein -1 multiply karo aur 35 dusri side le jao

OpenStudy (jiteshmeghwal9):

pehle -1 multiplly karu yaa 35 dudri side le jaaun ?

Parth (parthkohli):

Kuch bhi kar lo yaar!

OpenStudy (jiteshmeghwal9):

-1(35+5yx-24x)=0 -35-5yx+24x=0 -5yx+24x=35

Parth (parthkohli):

haan, LHS factor karo

OpenStudy (jiteshmeghwal9):

-x(5y-24)

Parth (parthkohli):

Yes, ab dhyan se dekho:\[35 = 1\times 35 \\ 35 = 7\times 5 \\ 35 = 5\times 7 \\ 35 = 35\times 1 \]

Parth (parthkohli):

Toh dekho, hum dekh sakte hain ki \(-x\) aur \(5y - 24\) ka product \(35\) hai, right?

OpenStudy (jiteshmeghwal9):

yes!

Parth (parthkohli):

Par jo maine upar post kiya, jaise \(1,35\) aur \(7,5\) etc., unka bhi toh product \(35\) hai na!

OpenStudy (jiteshmeghwal9):

haan

Parth (parthkohli):

Matlab ki humara solution inmein se ek equations se niklega:\[-x = 1, \ \ 5y - 24 =35 \\ -x = 7, \ 5y-24 =5 \\ -x = 5, \ 5y - 24 = 7 \\ -x = 35, 5y - 24 = 1\]

Parth (parthkohli):

Jis bhi equation mein dono \(x\) aur \(y\) integer ho, wohi answer hai!

OpenStudy (jiteshmeghwal9):

4th equation :- x=-35 y=5

Parth (parthkohli):

Yes =) ab answer check karo

OpenStudy (jiteshmeghwal9):

book ke answer se toh yeh match nahi kar raha hai

Parth (parthkohli):

Arrey book ko chhodo, book khoosat hoti hain ;)

OpenStudy (jiteshmeghwal9):

haha, ok i gt it, thanx :)

Parth (parthkohli):

You're welcome ^_^